OpenStudy (anonymous):

I need help with this calculus question. Some ancient Egyptians decided to build a pyramid with a square base that measured 750 feet on a side. The height of the pyramid was 500 feet. The density of the rock used in construction was 120 pounds per cubic foot. It was assumed that the work force would remain constant. Each laborer could do 160 foot pounds of work per hour, working 12 hour days for 330 days per year for 20 years. Determine, under these conditions, the number of laborers needed to construct the pyramid.

OpenStudy (amistre64):

Is it a solid pyramid? you know, for ceremonial purposes :)

OpenStudy (anonymous):

(1/3)*(750)^2*500 ft^3 * 120lbs/ft^3 * 1hr/(160lbs)/12hrs

OpenStudy (anonymous):

yeah its solid

OpenStudy (anonymous):

etc

OpenStudy (anonymous):

i know its a definite integral equation that is the density * volume * distance(or ht.) I just can't figure out how to represent the volume.

OpenStudy (anonymous):

888 laborers

OpenStudy (amistre64):

(S) A(y) dy {0,h} or something like that.... but the formula for any conelike volume is (1/3)(Base Area)(height)...

OpenStudy (amistre64):

intuition would say, find the amount of rock you need, figure out how much one laborer does in the alloted time span, then multiply that by how many laborers it take to do the job..... but other than that, I dont know...

OpenStudy (anonymous):

\[\int\limits_{0}^{50}(120)(volume)(y)\] the equation to find the work needed is something like this, i just can't figure out the section for volume.

OpenStudy (anonymous):

50 should be 500

OpenStudy (amistre64):

volume is integrated thru taking sections of the base area as it moves thru the height....

OpenStudy (anonymous):

also the volume is representing the volume of a slice

OpenStudy (amistre64):

each slice is x^2 dy in volume right?

OpenStudy (amistre64):

and x changes with repect to height...or y

OpenStudy (amistre64):

what is the slope of the face?

OpenStudy (anonymous):

yeah basically, but the size of the slice changes since it is getting smaller going up

OpenStudy (anonymous):

i don't know that isn't given

OpenStudy (amistre64):

500/(750/2) would be a cross section of the "triangle"

OpenStudy (amistre64):

the slope of a face is 500/375....might need to reduce. the equation relating that to x and y is y = -500/375(x) + 500

OpenStudy (amistre64):

slope is 4/3 when reduced

OpenStudy (amistre64):

x increase by 2 for every 3/4 increase in y .... if that helps

OpenStudy (amistre64):

(3/2) (S) y dy {0,500} = (3/4) y^2 {0,500} 500^2 = 250000(3/4) = 750000/4 = 187500 that matches 500(750)/2 = 187500 for a flat area..... not sure how it helps tho :)

OpenStudy (amistre64):

x^2 is the area of a normal base and it changes with respect to "y" at a rate of (3/2)(y) dy ?? It looks right, but this is still shakey ground fer me :)

OpenStudy (anonymous):

yeah i think ive gotten it now, thanks.

OpenStudy (amistre64):

if I messed up anything, youlet me know :)